OpenStudy (anonymous):

the width W of a rectanglar swimming pool is x+4 the area A is 2x^3-29x+12. What is an expression for the length of the pool?

OpenStudy (johnweldon1993):

Well area...= Length times Width So 2x^3 - 29x + 12 = Length * x + 4 so solve for Length...divide both sides by (x + 4) 2x^3 - 29x + 12 = Length * x + 4 ------------------------------ x + 4 x + 4 \[Length = \frac{ 2x^3 - 29x + 12 }{ x + 4 }\] Can you solve this...?